differentiate the following function with respect to x
log(x^2+x+1/x^2-x+1)

Grade:12

2 Answers

Rinkoo Gupta
9 years ago
y=log(x2+x+1)-log(x2-x+1)
y=log{x-1)(x2+x+1)/(x-1)} -log{((x+1)(x2-x+1)/(x+1)}
=log(x3-1)-log((x-1) -log(x3+1) +log(x+1)
=log((x3-1)/(x3+1)) +log((x+1)/(x-1))

Thanks & Regards
Rinkoo Gupta
AskIITians Faculty
Arun Kumar IIT Delhi
9 years ago
Hi
Please evaluate yourself.

Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty